(x^2-2x-15)(x-4)=0 solve by finding roots

Answer by ptaylor(2198)

(x^2-2x-15)(x-4)=0 solve by finding roots
Note that x^2-2x-15 has factors of (x-5)(x+3). How did I know this?? I determined that it has factors by inspection and by knowing that when a quadratic (Ax^2+Bx+C=0) has the A coefficient equal to 1, then the B coefficient is equal to the sum of the factors of the C coefficient, if it is factorable and in this case (-5)(+3) works.

So we have:
(x-5)(x+3)(x-4)=0
x=5
x=-3
x=4

Hope this helps-------ptaylor